OpenStudy (anonymous):

What is the measure of ∠RST? A. 70° B. 75° C. 85° D. 105°

OpenStudy (anonymous):

You can't tell anything but that it's not 105.

OpenStudy (anonymous):

ok but answer

OpenStudy (welshfella):

You can only guess - looks about 75 degrees but i could be wrong

OpenStudy (anonymous):

ok but do you guys know

OpenStudy (welshfella):

you can only measure it accurately with a protractor which i havent got !!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Which equation can be used to correctly find the value of m? A. 81 – m = 125 B. 125 + m = 81 C. 125 – 81 = m D. 81 + 125 = m

OpenStudy (welshfella):

its not 85 which will be closer to 90. Its 70 or 75

OpenStudy (anonymous):

a.

OpenStudy (anonymous):

hello ?

OpenStudy (anonymous):

What is the measure of ∠1? m∠1 = ° m∠2 = 50° m∠3 = 80° Three angles labeled 1, 2, and 3 are defined by rays. Angles 2 and 3 are adjacent. The non-shared rays from angles 2 and 2 form angle 1.

OpenStudy (anonymous):

the total is 360 subtract 50 and 80 leaving you with 230 for angle 1

OpenStudy (anonymous):

ok so 230